# Evaluation of LLMs for mathematical problem solving

### v2

@@ Abstract
Large Language Models (LLMs) have shown impressive performance on a range of educational tasks, but are still understudied for their potential to solve mathematical problems. In this study, we compare three prominent LLMs, including GPT-4o, DeepSeek-V3, and Gemini-2.0, on three mathematics datasets of varying complexities (GSM8K, MATH500, and University datasets). We take a five-dimensional approach based on the Structured Chain-of-Thought (SCoT) framework to assess final answer correctness, step completeness, step validity, intermediate calculation accuracy, and problem comprehension. The results show that GPT-4o is the most stable and consistent in performance across all the datasets, but particularly it performs outstandingly in high-level questions of the UNSW dataset. DeepSeek-V3 is competitively strong in well-structured domains such as optimisation, but suffers from fluctuations in accuracy in statistical inference tasks. Gemini-2.0 shows strong linguistic understanding and clarity in well-structured problems but performs poorly in multi-step reasoning and symbolic logic. Our error analysis reveals particular deficits in each model: GPT-4o is at times lacking in sufficient explanation or precision; DeepSeek-V3 leaves out intermediate steps; and Gemini-2.0 is less flexible in mathematical reasoning in higher dimensions.
@@
Large language models, ChatGPT-4o, DeepSeek-V3, Gemini-2.0, SCoT, Mathematical problem solving
@@ 1 Introduction
Progress in natural language processing (NLP)  [16] based on artificial intelligence technologies [17] and the advent of large language models (LLMs)  [18, 19, 20] have led to a revolutionary change in educational technology, giving rise to novel possibilities for mathematics education. The application of LLMs  [21, 22, 23] has fundamentally changed the design of human-computer interaction in educational settings  [24, 25] and initiated a revolutionary change in instructional approaches. In contrast to conventional systems that rely on predefined rules, large language models (LLMs) demonstrate more advanced language comprehension and generation capabilities  [26]. Not only do they accurately comprehend user intent and produce natural, fluent text content  [19], but they also adapt their output dynamically to feedback, thus creating a more semantically rich and interactive learning experience  [27]. These characteristics create new opportunities for applying LLMs and it has significantly driven the evolution of human-machine collaboration models in multiple fields  [28], including educational tutoring, customer service, programming assistance, and interactive communication support, enabling them to achieve better knowledge acquisition and problem-solving in learning, work. Large language models have demonstrated significant potential in mathematics education. Students have widely utilised LLMs to obtain problem-solving [18] strategies and answers [18], significantly shortening the time required to brainstorm and solve problems. However, whether this has truly improved learning outcomes remains to be explored [29]. Additionally, LLMs have the ability to accurately analyse the student’s learning situations[30] and dynamically adjust the presentation of content based on their pace and preferences[30], which can help promote "anytime, anywhere learning" teaching methods[31]. As LLM technology continues to advance, the potential for its use in education is becoming more widespread. The deep integration of mathematics education with artificial intelligence [32] is opening up new learning pathways and driving the construction of efficient, personalised learning platforms, thereby improving teaching quality and efficiency and promoting the comprehensive development of students[33]. Modern LLMs are typically built on deep learning models [34, 35] and widely applied in NLP tasks [36][37]. These models demonstrate strong capabilities in text understanding and generation capabilities and are widely applied in machine translation, text summarisation, and code generation [38]. Their task versatility stems from large-scale pretraining [39] and their capacity to manage tasks in both zero-shot and few-shot learning [40]. LLMs include GPT-4o [41], Gemini-2.0[42], and DeepSeek-V3[43] capable of text generation, logic reasoning, and multilingual processing[44]. Despite their advantages, LLMs encounter several challenges, such as hallucination (i.e., generating false information[21], and inherent data biases reflecting social and cultural stereotypes [45][46][47][48]. Although current LLMs demonstrate outstanding capabilities in mathematical problem solving, formula derivation, and theorem explanation [49], they remain weak in symbolic computation, multi-step algorithms and formal theorem proving [50]. To address these shortcomings, researchers have developed tools such as Wolfram Alpha[51] to enhance computational accuracy and logic rigour, and have improved programming performance through models like GitHub Copilot[52], GPT, and Code Llama[53]. However, these models still face challenges such as a lack of logical consistency[54], difficulty handling large-scale dependency structures, and generating errors or fictional functions when dealing with complex tasks. Algorithmic breakthroughs are still required to meet the demands of complex application scenarios [55], and how to efficiently utilise LLMs to improve learning outcomes while ensuring fairness [56] has become a focal point of current academic research.
Therefore, it is necessary to conduct systematic evaluations of LLMs using mathematical benchmark [57] datasets such as GSM8k [58] and MATH, which are widely used in mathematical tasks. And in order to study the performance of LLMs in advanced-level mathematical problems, we plan to construct the University of New South Sales problem dataset, which is collected by using assignment and final exams from the Mathematics and Statistics School. These evaluations help us to identify the strengths and weakness of models across different tasks [59], and provide empirical support for model optimisation. Although existing studies have fully demonstrated the ability of LLMs in mathematical and programming tasks, their shortcomings in logical reasoning and computational accuracy[60] still limit their widespread application. Therefore, it is imperative to continue optimising LLMs from the perspectives of algorithm design and technical pathways to enhance their overall performance and meet the increasingly complex and diverse demands of practical applications[61].
Our study is based on this premise and aims to evaluate the performance of three mainstream LLMs, GPT-4o, DeepSeek-V3, and Gemini-2.0, in solving mathematical problems. We will examine their capabilities in terms of accuracy, reasoning quality, completeness, and consistency using a set of problems covering different levels of difficulty. We will use an automated quantitative evaluation framework to provide theoretical support and practical evidence for their application in education.
Our study will focus on the following contributions:
@@ 3.1 Mathematical problems
    University Graduate Question Bank: This dataset consists of exam and assignment questions covering probability inference, matrix decomposition, optimisation theory, and other courses. The language is rigorous, and the expression style is academic. Some questions have open structures, conceptual explanations, and non-standard expressions. The details about this dataset can be shown in Table LABEL:tab:data_unsw.
@@
University Problem 130
@@ 3.2 Automated LLM Evaluation Framework
    University Problem dataset (real academic assessment scenarios (Table LABEL:tab:data_unsw))
@@
    Stage 1 - Problem definition: We establish the foundations of the evaluation using the SCoT (Scoring Chain-of-Thought (SCoT) [115] framework that emphasises systematic evaluation of the completeness and logical consistency of the reasoning chain to test the performance of LLMs, which serves as the basis for the four subsequent evaluation indicators design.
@@ 4.1 LLMs for solving GSM8K problems
We investigated more about the errors contributing to this mediocre result, including interruptions in the reasoning chain, omissions of steps, misunderstandings of quantitative relationships, and errors in unit conversion. In the case of GSM8K, all three models sometimes arrives at the correct answer. However, they cannot generate intermediate steps due to an unstable reasoning process or improper handling of details, which reflects their lack of coherence and robustness in multi-step tasks, as shown in Figure LABEL:fig:sample_question_answer.

### v3

@@ Abstract
Large Language Models (LLMs) have shown impressive performance on a range of educational tasks, but are still understudied for their potential to solve mathematical problems. In this study, we compare three prominent LLMs, including GPT-4o, DeepSeek-V3, and Gemini-2.0, on three mathematics datasets of varying complexities (GSM8K, MATH500, and University datasets). We take a five-dimensional approach based on the Structured Chain-of-Thought (SCoT) framework to assess final answer correctness, step completeness, step validity, intermediate calculation accuracy, and problem comprehension. The results show that GPT-4o is the most stable and consistent in performance across all the datasets, but particularly it performs outstandingly in high-level questions of the University dataset. DeepSeek-V3 is competitively strong in well-structured domains such as optimisation, but suffers from fluctuations in accuracy in statistical inference tasks. Gemini-2.0 shows strong linguistic understanding and clarity in well-structured problems but performs poorly in multi-step reasoning and symbolic logic. Our error analysis reveals particular deficits in each model: GPT-4o is at times lacking in sufficient explanation or precision; DeepSeek-V3 leaves out intermediate steps; and Gemini-2.0 is less flexible in mathematical reasoning in higher dimensions.
@@
Large language models, GPT-4o, DeepSeek-V3, Gemini-2.0, SCoT, Mathematical problem solving
@@ 1 Introduction
Progress in natural language processing (NLP)  [16] based on artificial intelligence technologies [17] and the advent of large language models (LLMs)  [18, 19, 20] have led to a revolutionary change in educational technology, giving rise to novel possibilities for mathematics education. The application of LLMs  [21, 22, 23] has fundamentally changed the design of human-computer interaction in educational settings  [24, 25] and initiated a revolutionary change in instructional approaches. In contrast to conventional systems that rely on predefined rules, large language models (LLMs) demonstrate more advanced language comprehension and generation capabilities  [26]. Not only do they accurately comprehend user intent and produce natural, fluent text content  [19], but they also adapt their output dynamically to feedback, thus creating a more semantically rich and interactive learning experience  [27]. These characteristics create new opportunities for applying LLMs and it has significantly driven the evolution of human-machine collaboration models in multiple fields  [28], including educational tutoring, customer service, programming assistance, and interactive communication support, enabling them to achieve better knowledge acquisition and problem-solving in learning, work. Large language models have demonstrated significant potential in mathematics education. Students have widely utilised LLMs to obtain problem-solving [18] strategies and answers [18], significantly shortening the time required to brainstorm and solve problems. However, whether this has truly improved learning outcomes remains to be explored [29]. Additionally, LLMs have the ability to accurately analyse the student’s learning situations[30] and dynamically adjust the presentation of content based on their pace and preferences[30], which can help promote "anytime, anywhere learning" teaching methods[31]. As LLM technology continues to advance, the potential for its use in education is becoming more widespread. The deep integration of mathematics education with artificial intelligence [32] is opening up new learning pathways and driving the construction of efficient, personalised learning platforms, thereby improving teaching quality and efficiency and promoting the comprehensive development of students[33]. Modern LLMs are typically built on deep learning models [34, 35] and widely applied in NLP tasks [36][37]. These models demonstrate strong capabilities in text understanding and generation capabilities and are widely applied in machine translation, text summarisation, and code generation [38]. Their task versatility stems from large-scale pretraining [39] and their capacity to manage tasks in both zero-shot and few-shot learning [40]. LLMs include GPT-4o [41], Gemini-2.0[42], and DeepSeek-V3[43] capable of text generation, logic reasoning, and multilingual processing[44]. Despite their advantages, LLMs encounter several challenges, such as hallucination (i.e., generating false information[21], and inherent data biases reflecting social and cultural stereotypes [45][46][47][48]. Although current LLMs demonstrate outstanding capabilities in mathematical problem solving, formula derivation, and theorem explanation [49], they remain weak in symbolic computation, multi-step algorithms and formal theorem proving [50]. To address these shortcomings, researchers have developed tools such as Wolfram Alpha[51] to enhance computational accuracy and logic rigour, and have improved programming performance through models like GitHub Copilot[52], GPT, and Code Llama[53]. However, these models still face challenges such as a lack of logical consistency[54], difficulty handling large-scale dependency structures, and generating errors or fictional functions when dealing with complex tasks. Algorithmic breakthroughs are still required to meet the demands of complex application scenarios [55], and how to efficiently utilise LLMs to improve learning outcomes while ensuring fairness [56] has become a focal point of current academic research. Therefore, it is necessary to conduct systematic evaluations of LLMs using mathematical benchmark [57] datasets such as GSM8k [58] and MATH, which are widely used in mathematical tasks. And in order to study the performance of LLMs in advanced-level mathematical problems, we plan to construct the University of New South Sales problem dataset, which is collected by using assignment and final exams from the Mathematics and Statistics School. These evaluations help us to identify the strengths and weakness of models across different tasks [59], and provide empirical support for model optimisation. Although existing studies have fully demonstrated the ability of LLMs in mathematical and programming tasks, their shortcomings in logical reasoning and computational accuracy[60] still limit their widespread application. Therefore, it is imperative to continue optimising LLMs from the perspectives of algorithm design and technical pathways to enhance their overall performance and meet the increasingly complex and diverse demands of practical applications[61]. Our study is based on this premise and aims to evaluate the performance of three mainstream LLMs, GPT-4o, DeepSeek-V3, and Gemini-2.0, in solving mathematical problems. We will examine their capabilities in terms of accuracy, reasoning quality, completeness, and consistency using a set of problems covering different levels of difficulty. We will use an automated quantitative evaluation framework to provide theoretical support and practical evidence for their application in education. Our study will focus on the following contributions:
@@ 3.1 Mathematical problems
    University Graduate Question Bank: This dataset consists of exam and assignment questions covering probability inference, matrix decomposition, optimisation theory, and financial computation, which are from the MIT Open Courseware. The language is rigorous, and the expression style is academic. Some questions have open structures, conceptual explanations, and non-standard expressions. The details about this dataset can be shown in Table 1.
    Table 1: Summary of mathematical problem dataset from MIT Open Courseware.
    Subject Code Title Assessment Type Year
    Optimisation 16.323 Principles of Optimal Control Final Exam 2007
    15.093 Optimisation Methods Final Exam 2003/2006/2008
    6.253 Convex Analysis and Optimisation Midterm Exam 2010/2012
    18085 Computational Science and Engineering Exam 2008
    Financial Computational Math 18.330 Introduction to Numerical Analysis Homework 2012
    18.336 Numerical Methods of Applied Mathematics Problem Set 2009
    Statistical Inference 18.440 Probability and Random Variables Final Exam 2014
    18.655 Mathematical Statistics Midterm Exam 2016
    15.085J Fundamentals of Probability Final Exam 2018
    6.431 Probabilistic System Analysis Final Exam 2006
@@
University Problem 140 Advanced MIT Open Courseware
@@ 3.2 Automated LLM Evaluation Framework
    University Problem dataset (real academic assessment scenarios
@@
    Stage 1 - Problem definition: We establish the foundations of the evaluation using the SCoT (Structure Chain-of-Thought (SCoT) [115] framework that emphasises systematic evaluation of the completeness and logical consistency of the reasoning chain to test the performance of LLMs, which serves as the basis for the five subsequent evaluation indicators design.
@@
    University A quantitative analysis was conducted on 140 mathematical problems from various subjects from the MIT Open Courseware to study the performance of language models in advanced mathematics.
@@ 4.1 LLMs for solving GSM8K problems
We investigated more about the errors contributing to this mediocre result, including interruptions in the reasoning chain, omissions of steps, misunderstandings of quantitative relationships, and errors in unit conversion. In the case of GSM8K, all three models sometimes arrives at the correct answer. However, they cannot generate intermediate steps due to an unstable reasoning process or improper handling of details, which reflects their lack of coherence and robustness in multi-step tasks.